I'm in college. I don't understand how to get the variables.
Here is my problem
12x + 3 = 8x - 24
Can you show me step by step with no short cuts how to get this answer.

Answer by confusedgirl81(14)

First thing is to get like terms together

12x + 3 = 8x - 24 {subtract 8x from both sides}
12x - 8x + 3 = 8x - 8x - 24 {combine like terms}
4x + 3 = -24 {subtract 3 from both sides}
4x = -24 - 3
4x = -27 {divide both sides by 4}
x = -27/4


So x = -27/4


If you need more help, just ask!